# references.bib

### v1

@@ 2.2 Structure of spectral extremal graphs
We now prove Theorem 2. Let $\mathcal{F}$ be as in Theorem 2 and let $k=k(\mathcal{F})$ as in 1. If $k=1$, then the conclusion of Theorem 2 follows from Theorem 1, so we will assume that $k\geq 2$ whenever it is convenient. Similarly to the previous subsection, we have the following lemma.
@@
Let $m$ be a large enough even constant satisfying 16. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Suppose $G\in\mathrm{SPEX}(n,\mathcal{F})$, where $n$ is large enough. Choose positive constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$,$\varepsilon_{4},\varepsilon_{5},\varepsilon_{6}$ satisfying the following inequalities (we do not attempt to avoid redundancy).
(i) $2(r+1)\sqrt{\varepsilon}<\min\{\varepsilon_{1},\varepsilon_{3}\}$ (ii) $2\varepsilon_{1}+\sqrt{\varepsilon}<1/r$ (iii) $2(r+1)^{2}(\sqrt{\varepsilon}+\varepsilon_{1})<\varepsilon_{2}$ (iv) $M\sqrt{\varepsilon}<1/r$ (v) $\sqrt{\varepsilon}+\varepsilon_{2}+2\varepsilon_{3}<1/r$ (vi) $\sqrt{\varepsilon}+\varepsilon_{2}+\varepsilon_{3}<\varepsilon_{4}$ (vii) $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{3}$ (viii) $\max\{\varepsilon_{4}+\sqrt{\varepsilon},3\sqrt{\varepsilon}\}<\varepsilon_{5}$ (ix) $\sqrt{\varepsilon}+m(k+1)\varepsilon_{5}<1/r$ (x) $\sqrt{\varepsilon}+m(k+r)\varepsilon_{5}<1/r$ (xi) $\sqrt{\varepsilon}+M\varepsilon_{5}<1/r$ (xii) $\frac{r(\varepsilon_{1}+(r-1)\varepsilon_{5})}{r-1}<\varepsilon_{6}$ (xiii) $\frac{r}{r-1}(\varepsilon_{5}+\varepsilon_{1}+3\sqrt{\varepsilon}+\varepsilon_{2}/r+(r-1)\varepsilon_{5})<\varepsilon_{6}$ (xiv) $(m+1)\varepsilon_{6}<1$ (xv) $\sqrt{\varepsilon}+2mr\varepsilon_{5}<1/r$ (xvi) $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$ (xvii) $2\varepsilon_{6}<1-\frac{k-2}{k-1}$.
@@
since $2(r+1)\sqrt{\varepsilon}<\varepsilon_{1}$, (i). Thus, we eventually find a copy of $K_{1,m}+T_{m(r-1),r-1}$ in $G$, contradicting 16. So some such $i_{1}$ exists. To prove the uniqueness, suppose there are $i\neq j$ such that $d_{V_{i}-X}(z),d_{V_{j}-X}(z)\leq\varepsilon_{1}n+O(1)$; then we have $d(z)\leq\left(1-\frac{2}{r}+2\varepsilon_{1}\right)n+O(1)$ which is a contradiction by (ii). Hence, there is a unique $i$ such that $d_{V_{i}-X}(z)\leq\varepsilon_{1}n$. Finally, since $d(z)\geq\left(1-\frac{1}{r}\right)n+O(1)$, we have that for any $j\neq i$, we have $d_{V_{j}}(z)\geq(1/r-\varepsilon_{1})n+O(1)$. ∎
@@
$=\lambda-\left(2(r+1)\sqrt{\varepsilon}+\varepsilon_{1}+2\varepsilon_{1}\right)n+O(1)$
@@
The next section of the proof follows the lines of the argument in subsection 2.1. Define an ordinary vertex $v\in V_{i}$ to be one such that $N(v)=\bigcup_{j\neq i}V_{j}-X$.
@@
If $G^{\prime}$ is obtained from $G$ by deleting some vertex and adding an ordinary vertex $u\in V_{i}$, then $G^{\prime}$ is $\mathcal{F}$-free.
@@
since $\sqrt{\varepsilon}+\varepsilon_{2}+\varepsilon_{3}<\varepsilon_{4}$, (vi), which by 8 contradicts $G\in\mathrm{SPEX}(n,\mathcal{F})$. Finally, suppose there is no index $i$ for which $d_{V_{i}}(v)\leq\varepsilon_{3}n$. Then using $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{3}$, (vii) and a similar argument to subsection 2.1, we find a $K_{1,m}+T_{m(r-1),r-1}$, which contradicts 16. ∎
By 21, we may define a new partition $V=W_{1}\sqcup\cdots\sqcup W_{r}$ as follows. For $v\in V_{i}-X$, let $v\in W_{i}$. If $v\in X$, then let $v\in W_{i}$ where $i$ is chosen so that $d_{V_{i}}(v)\leq\varepsilon_{3}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{4}n$ for each $j\neq i$. Let $w_{i}:=|W_{i}|$. Note that $w_{i}\in(n/r-\sqrt{\varepsilon}n+O(1),n/r+\sqrt{\varepsilon}n+O(1))$. Thus for each $v\in W_{i}$, for each $j\neq i$, either $d_{W_{j}}(v)\geq w_{j}-(\varepsilon_{4}+\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{5}n$ or $d_{W_{j}}(v)\geq w_{j}-(2\sqrt{\varepsilon}+\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{5}n$, since $\max\{\varepsilon_{4}+\sqrt{\varepsilon},3\sqrt{\varepsilon}\}<\varepsilon_{5}$, (viii). Also, redefine an ordinary vertex $v\in W_{i}$ to mean one for which $N(v)=V-W_{i}$.
@@
Let $G^{\prime}$ be obtained from $G$ by deleting some vertex and adding an ordinary vertex $u\in V_{i}$. Then $G^{\prime}$ is $\mathcal{F}$-free.
@@
This follows immediately from 22, 23, and 16. ∎
@@
For each $v\in W_{i}$, for each $j\neq i$, we have $d_{W_{j}}(v)\geq w_{i}-m$.
@@
Suppose not. By 23, the graph $G^{\prime}$ obtained by replacing $v$ with an ordinary vertex is $\mathcal{F}$-free, and by 24 and 25 we have
@@
Assume without loss of generality that $W_{1}$ is the part for which $e(W_{i})\to\infty$ is possible, so that $e(W_{i})=O(1)$ for all $i\neq 1$. Observe that for each $i\neq 1$, we may write $W_{i}=W_{i}^{o}\sqcup W_{i}^{m}$, where $e(W_{i}^{o})=0$ and $|W_{i}^{m}|=O(1)$. We will also define $W^{o}=\bigcup_{i=1}^{r}W_{i}^{o}$ and $W^{m}=\bigcup_{i=1}^{r}W_{i}^{m}$.
@@
By 25, the change in $\frac{1}{2}x^{T}Ax$ is at least
$b(1-\varepsilon_{6})^{2}-a>b(1-2\varepsilon_{6})-a>0.$
@@
We can write $W_{1}=W_{1}^{o}\sqcup W_{1}^{m}$, where $W_{1}^{o}$ is a union of trees of fixed order $k$, each tree component of $G[W_{1}^{o}]$ occurs $\omega(1)$ times, and $|W_{1}^{m}|=O(1)$.
@@
The graph $G[W_{1}]$ has at most $m$ components whose order is at least $k+1$.
@@
Similarly to 6, we find an $\mathcal{F}$-free graph $G^{\prime}$ obtained from $G$ by adding $b=w_{1}\frac{k-1}{k}+o(n)$ edges and deleting $a=w_{1}\frac{k-2}{k-1}+o(n)$ edges. Then
@@
As in 7, we obtain an $\mathcal{F}$-free graph $G^{\prime}$ obtained from $G$ by deleting $a=k(\ell-1)$ edges and adding $b=k\ell$ edges, where $\ell\leq k-1$. Then we have
$\frac{a}{b}\leq\frac{\ell-1}{\ell}\leq\frac{k-2}{k-1}<1-2\varepsilon_{6}$
since $2\varepsilon_{6}<1-\frac{k-2}{k-1}$, (xvii). Hence, 29 gives $\lambda(G^{\prime})>\lambda(G)$, contradicting $G\in\mathrm{SPEX}(n,\mathcal{F})$. ∎
Combining 9, 10, and 13, we obtain the result (we put any trees of order $k$ which occur only $O(1)$ times in $W_{1}^{m}$). ∎
@@
Now we know that $G$ can be obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees of order $k$ in $W_{1}$, then adding and deleting $O(1)$ edges. Now we determine the values of the $w_{i}$. Since all vertices in $W_{i}^{o}$ ($i\neq 1$) have the same neighborhood, they have the same eigenweight, so let $x_{i}$ be the common eigenweight of vertices in $W_{i}^{o}$. Let $x_{1}=\frac{1}{\lambda}\sum_{v\not\in W_{1}}x_{v}$. Observe that $\lambda=\Theta(n)$, a fact we will use extensively below.
@@
$\lambda x_{j}$ $=\lambda x_{i}+x_{i}(w_{i}-|W_{i}^{m}|)+\sum_{w\in W_{i}^{m}}x_{w}-w_{j}(w_{j}-|W_{j}^{m}|)-\sum_{w\in W_{j}^{m}}x_{w}$
$=\lambda x_{i}+w_{i}x_{i}+|W_{i}^{m}|O(n^{-1})-w_{j}x_{j}-W_{j}^{m}O(n^{-1})$
@@
The spectral version of 13 gives the following fact: for any $u\in W_{i}^{o}$, for any $j\neq i$, we have $\lambda(G_{ij})\leq\lambda(G)$ and so
@@
$X_{1}=\sum_{u\in W_{1}}(x_{1}+\frac{d_{W_{1}}(u)}{\lambda}x_{1}+O(n^{-2}))=w_{1}x_{1}+\frac{2e(W_{1})}{\lambda}x_{1}+O(n^{-1}).$
@@
00 0 $\geq w_{1}x_{1}+\frac{2e(W_{1})}{\lambda}x_{1}-w_{i}x_{i}+O(1)=w_{1}x_{1}-w_{i}(x_{1}+O(n^{-1}))+O(1)$
$O(1)$ $\geq w_{1}-w_{i}$
@@
From the bounds proved so far it is not too difficult to show that $\lambda=\frac{r-1}{r}n+O(1)$, which will be useful for proving the estimates below.
@@
Now since $w_{i}<w_{j}$, we have $w_{i}<n/2$, and since $r\geq 3$ we have $\lambda\geq\left(1-\frac{1}{r}\right)n+O(1)\geq 3n/2+O(1)$. This implies that $\lambda>w_{i}+2n^{1/2}$, which together with $w_{j}\geq w_{i}+2$ implies $\lambda(w_{j}-w_{i}-1)>w_{i}+2n^{1/2}$, so $(w_{j}-1)\frac{\lambda+w_{i}}{\lambda+w_{j}}-w_{i}\geq\Omega(n^{-1/2})$ and so the expression displayed above is positive, a contradiction. ∎
@@
$\geq 3\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)-2+O(n^{-1})$
@@
Now using $e(W_{1})\leq\frac{n}{r}+O(1)$ we obtain $\frac{2e(W_{1})}{\lambda}\left(\frac{w_{i}}{\lambda+w_{i}}-1\right)\geq-\frac{2}{r}+O(n^{-1})$, so
$00\geq 3\left(1-\frac{1}{r}\right)-1-\frac{2}{r}+O(n^{-1})\geq 2-\frac{5}{3}+O(n^{-1})>$ 0
@@
Let $x$ be the principal eigenvector of $H$. From the estimates proved above, we already have that $x_{v}=1+O(n^{-1})$ for every $v$. Let $A^{\prime}$ be the set of deleted edges and $B^{\prime}$ be the set of added edges. Then we have
@@
which gives a contradiction if $r\geq 4$. If $r=3$, we require a more detailed argument. Without loss of generality, $i=2$. Let $G^{\prime}$ be obtained from $G$ by converting every vertex in $W^{m}$ to an ordinary vertex, let $a=e(G^{\prime})-e(G)=O(1)$, and let $G_{21}^{\prime}=(G^{\prime})_{21}$. If $a=E(W^{m})$, then observe by 41 that $\lambda(G^{\prime})=\lambda(G)-2a/n+O(n^{-2})$ and $\lambda(G_{21})=\lambda(G_{21}^{\prime})+2a/n+O(n^{-2})$, and so it suffices to show that $\lambda(G_{21}^{\prime})-\lambda(G^{\prime})\geq\delta/n$ for some constant $\delta>0$. Let $V(G_{21}^{\prime})$ have partition $U_{1}\sqcup U_{2}\sqcup U_{3}$ so that $|U_{1}|=w_{1}+1$, $|U_{2}|=w_{2}-1$, and $|U_{3}|=w_{3}$. Note that according to the previous lemmas, we either have $w_{1}=w_{2}-2=w_{3}-1$ or $w_{1}=w_{2}-2=w_{3}-2$. Let $\mu=\lambda(G^{\prime})$, $\rho=\lambda(G_{21}^{\prime})$, with corresponding eigenvectors $y,z$, respectively, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}x_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in U_{1}}z_{v}=1$. Let $y_{2},y_{3}$ be the common eigenweights of vertices in $W_{2},W_{3}$, and let $z_{2},z_{3}$ be the common eigenweights of vertices in $U_{2},U_{3}$.
@@
$\lambda(G_{21}^{\prime})-\lambda(G^{\prime})\geq\frac{-w_{1}(1+O(n^{-1}))+w_{2}(1+O(n^{-1}))}{y^{T}y}=\frac{O(1)}{y^{T}y}=O(n^{-1}).$
@@
Let $Y_{1}=\sum_{v\in W_{1}}y_{v}$, $Z_{1}=\sum_{v\in U_{1}}z_{v}$. Then using the eigenweight estimates above we obtain
@@
Suppose $G_{1}$, $G_{2}$ are two graphs obtained from $T_{n,r}$ by embedding the same number of trees of order $k\geq 4$ in a large part so as to cover all but $O(1)$ vertices in the part. Then
@@
Let $G_{\mathrm{star}}$ and $G_{\mathrm{path}}$ be the graphs obtained from $G_{1}$ or $G_{2}$ be replacing each embedded tree with a star or path, respectively.

### v2

@@ 2.2 Structure of spectral extremal graphs
We now prove Theorem 2. Let $\mathcal{F}$ be as in Theorem 2 and let $k=k(\mathcal{F})$ as in 1. If $k=1$, then the conclusion of Theorem 2 follows from Theorem 1, so we will assume that $k\geq 2$. Similarly to the previous subsection, we have the following lemma.
@@
Let $m$ be a large enough even constant satisfying 16. Let $M=\max\{|V(F)|:F\in\mathcal{F}\}$. Suppose $G\in\mathrm{SPEX}(n,\mathcal{F})$, where $n$ is large enough. Choose positive constants $\varepsilon,\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$,$\varepsilon_{4},\varepsilon_{5},\varepsilon_{6}$ satisfying the following inequalities (we do not attempt to avoid redundancy or optimize the constants).
(i) $2(r+1)\sqrt{\varepsilon}<\min\{\varepsilon_{1},\varepsilon_{3}\}$ (ii) $2\varepsilon_{1}+\sqrt{\varepsilon}<1/r$ (iii) $2(r+1)^{2}(\sqrt{\varepsilon}+\varepsilon_{1})<\varepsilon_{2}$ (iv) $M\sqrt{\varepsilon}<1/r$ (v) $\sqrt{\varepsilon}+\varepsilon_{2}+2\varepsilon_{3}<1/r$ (vi) $\sqrt{\varepsilon}+\varepsilon_{2}+\varepsilon_{3}<\varepsilon_{4}$ (vii) $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{3}$ (viii) $\max\{\varepsilon_{4}+\sqrt{\varepsilon},3\sqrt{\varepsilon}\}<\varepsilon_{5}$ (ix) $\sqrt{\varepsilon}+m(k+1)\varepsilon_{5}<1/r$ (x) $\sqrt{\varepsilon}+m(k+r)\varepsilon_{5}<1/r$ (xi) $\sqrt{\varepsilon}+M\varepsilon_{5}<1/r$ (xii) $\frac{r(\varepsilon_{1}+(r-1)\varepsilon_{5})}{r-1}<\varepsilon_{6}$ (xiii) $\frac{r}{r-1}(\varepsilon_{5}+\varepsilon_{1}+3\sqrt{\varepsilon}+\varepsilon_{2}/r+(r-1)\varepsilon_{5})<\varepsilon_{6}$ (xiv) $(m+1)\varepsilon_{6}<1$ (xv) $\sqrt{\varepsilon}+2mr\varepsilon_{5}<1/r$ (xvi) $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$ (xvii) $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$.
@@
since $2(r+1)\sqrt{\varepsilon}<\varepsilon_{1}$, (i). Thus, we eventually find a copy of $K_{1,m}+T_{m(r-1),r-1}$ in $G$, contradicting 16. So some such $i_{1}$ exists. To prove the uniqueness, suppose there are $i\neq j$ such that $d_{V_{i}-X}(z),d_{V_{j}-X}(z)\leq\varepsilon_{1}n$; then we have $d(z)\leq\left(1-\frac{2}{r}+2\varepsilon_{1}\right)n+O(1)$ which is a contradiction by (ii). Hence, there is a unique $i$ such that $d_{V_{i}-X}(z)\leq\varepsilon_{1}n$. Finally, since $d(z)\geq\left(1-\frac{1}{r}\right)n+O(1)$, we have that for any $j\neq i$, we have $d_{V_{j}}(z)\geq(1/r-\varepsilon_{1})n+O(1)$. ∎
@@
$\geq\lambda-\left(2(r+1)\sqrt{\varepsilon}+\varepsilon_{1}+2r\varepsilon_{1}\right)n+O(1)$
@@
The next section of the proof follows the lines of the argument in subsection 2.1. Define a $V$-ordinary vertex $v\in V_{i}$ to be one such that $N(v)=\bigcup_{j\neq i}V_{j}-X$.
@@
If $G^{\prime}$ is obtained from $G$ by deleting some vertex and adding a $V$-ordinary vertex $u\in V_{i}$, then $G^{\prime}$ is $\mathcal{F}$-free.
@@
since $\sqrt{\varepsilon}+\varepsilon_{2}+\varepsilon_{3}<\varepsilon_{4}$, (vi), which by 8 contradicts $G\in\mathrm{SPEX}(n,\mathcal{F})$. Finally, suppose there is no index $i$ for which $d_{V_{i}}(v)\leq\varepsilon_{3}n$. Then using $(2mr+1)\sqrt{\varepsilon}<\varepsilon_{3}$, (vii) and a similar argument to 4, we find a $K_{1,m}+T_{m(r-1),r-1}$, which contradicts 16. ∎
By 21, we may define a new partition $V=W_{1}\sqcup\cdots\sqcup W_{r}$ as follows. For $v\in V_{i}-X$, let $v\in W_{i}$. If $v\in X$, then let $v\in W_{i}$ where $i$ is chosen so that $d_{V_{i}}(v)\leq\varepsilon_{3}n$ and $d_{V_{j}}(v)\geq n/r-\varepsilon_{4}n$ for each $j\neq i$. Let $w_{i}:=|W_{i}|$. Note that $w_{i}\in(n/r-\sqrt{\varepsilon}n+O(1),n/r+\sqrt{\varepsilon}n+O(1))$. Thus for each $v\in W_{i}$, for each $j\neq i$, either $d_{W_{j}}(v)\geq w_{j}-(\varepsilon_{4}+\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{5}n$ or $d_{W_{j}}(v)\geq w_{j}-(2\sqrt{\varepsilon}+\sqrt{\varepsilon})n+O(1)\geq w_{j}-\varepsilon_{5}n$, since $\max\{\varepsilon_{4}+\sqrt{\varepsilon},3\sqrt{\varepsilon}\}<\varepsilon_{5}$, (viii). Also, call a vertex $v\in W_{i}$ $W$-ordinary if $N(v)=V-W_{i}$.
@@
Let $G^{\prime}$ be obtained from $G$ by deleting some vertex and adding a $W$-ordinary vertex $u\in V_{i}$. Then $G^{\prime}$ is $\mathcal{F}$-free.
@@
This follows immediately from 22 and 16. ∎
@@
For each $v\in W_{i}$, for each $j\neq i$, we have $d_{W_{j}}(v)\geq w_{j}-m$.
@@
Suppose not. By 23, the graph $G^{\prime}$ obtained by replacing $v$ with a $W$-ordinary vertex is $\mathcal{F}$-free, and by 24 and 25 we have
@@
Assume without loss of generality that $W_{1}$ is the part for which $e(W_{i})\geq C$ is possible, so that $e(W_{i})=O(1)$ for all $i\neq 1$. Observe that for each $i\neq 1$, we may write $W_{i}=W_{i}^{o}\sqcup W_{i}^{m}$, where $e(W_{i}^{o})=0$ and $|W_{i}^{m}|=O(1)$. We will also define $W^{o}=\bigcup_{i=1}^{r}W_{i}^{o}$ and $W^{m}=\bigcup_{i=1}^{r}W_{i}^{m}$.
@@
By 25, we have
$\frac{1}{2}(x^{T}A(G^{\prime})x-x^{T}A(G)x)\geq b(1-\varepsilon_{6})^{2}-a>b(1-2\varepsilon_{6})-a>0.$
@@
We can write $W_{1}=W_{1}^{o}\sqcup W_{1}^{m}$, where $W_{1}^{o}$ is a union of trees of fixed order $k$, each tree component of $G[W_{1}^{o}]$ occurs more than $M(m+1)$ times, and $|W_{1}^{m}|=O(1)$.
@@
The number of components of $G[W_{1}]$ whose order is at least $k+1$ is at most $m$.
@@
Similarly to 5, we find an $\mathcal{F}$-free graph $G^{\prime}$ obtained from $G$ by adding $b=w_{1}\frac{k-1}{k}+o(n)$ edges and deleting $a=w_{1}\frac{k-2}{k-1}+o(n)$ edges. Then
@@
As in 7, we obtain an $\mathcal{F}$-free graph $G^{\prime}$ obtained from $G$ by deleting $a=k(\ell-1)$ edges and adding $b=(k-1)\ell$ edges, where $\ell\leq k-1$. Then we have
$\frac{a}{b}\leq\frac{k(\ell-1)}{(k-1)\ell}\leq\frac{k(k-2)}{(k-1)^{2}}<1-2\varepsilon_{6}$
since $2\varepsilon_{6}<1-\frac{k(k-2)}{(k-1)^{2}}$, (xvii). Hence, 29 gives $\lambda(G^{\prime})>\lambda(G)$, contradicting $G\in\mathrm{SPEX}(n,\mathcal{F})$. ∎
Combining 9, 10, and 13, we obtain the result (we put any trees of order $k$ which occur at most $M(m+1)$ times in $W_{1}^{m}$). ∎
@@
Now we know that $G$ can be obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees of order $k$ in $W_{1}$, then adding and deleting $O(1)$ edges. Now we determine the values of the $w_{i}$. Since all vertices in $W_{i}^{o}$ ($i\neq 1$) have the same neighborhood, they have the same eigenweight, so let $x_{i}$ be the common eigenweight of vertices in $W_{i}^{o}$. Let $x_{1}=\frac{1}{\lambda}\sum_{v\not\in W_{1}}x_{v}$. Observe that $\lambda=\Theta(n)$, a fact we will use frequently below.
@@
$\lambda x_{j}$ $=\lambda x_{i}+x_{i}(w_{i}-|W_{i}^{m}|)+\sum_{w\in W_{i}^{m}}x_{w}-x_{j}(w_{j}-|W_{j}^{m}|)-\sum_{w\in W_{j}^{m}}x_{w}$
$=\lambda x_{i}+w_{i}x_{i}+|W_{i}^{m}|O(n^{-1})-w_{j}x_{j}-|W_{j}^{m}|O(n^{-1})$
@@
The spectral version of 13 gives the following fact: for any $u\in W_{i}^{o}$, for any $j\neq i$, we have $\lambda(G_{ij})\leq\lambda(G)$ and so with $u$ being the transferred vertex, we have
@@
$X_{1}=\sum_{u\in W_{1}}(x_{1}+\frac{d_{W_{1}}(u)}{\lambda}x_{1}+O(n^{-2}))=w_{1}x_{1}+\frac{2e(W_{1})}{\lambda}x_{1}+O(n^{-1})=w_{1}x_{1}+O(1).$
@@
00 0 $\geq w_{1}x_{1}-w_{i}x_{i}+O(1)$
$=w_{1}x_{1}-w_{i}\frac{\lambda+w_{1}+O(1)}{\lambda+w_{i}}x_{1}+O(1)=\left(w_{1}-w_{i}\frac{\lambda+w_{i}}{\lambda+w_{1}}\right)x_{1}$
$O(n)$ $\geq(\lambda+w_{i})w_{1}-(\lambda+w_{1})w_{i}=\lambda(w_{1}-w_{i})$
@@
$=w_{i}\frac{\lambda+w_{1}+O(1)}{\lambda+w_{i}}x_{1}-w_{1}x_{1}+O(1)$
$=\left(w_{i}\frac{\lambda+w_{1}}{\lambda+w_{i}}-w_{1}\right)x_{1}+O(1)$
$O(n)$ $\geq(\lambda+w_{1})w_{i}-(\lambda+w_{i})w_{1}=\lambda(w_{i}-w_{1})$
@@
Combining 36 and 37 now gives $x_{i}=x_{1}+O(n^{-1})$ for every $i$. Furthermore, combining 37 with the bound $\lambda\leq\Delta(G)$ proves $\lambda\leq\frac{r-1}{r}n+O(1)$, hence $\lambda=\frac{r-1}{r}n+O(1)$. These estimates will be useful in the rest of the argument.
@@
Now since $w_{i}<w_{j}$, we have $w_{i}<n/2$, and since $r\geq 3$ we have $\lambda\geq\left(1-\frac{1}{r}\right)n+O(1)\geq 2n/3+O(1)$. This implies that $\lambda>w_{i}+n^{1/2}$, which together with $w_{j}\geq w_{i}+2$ implies $\lambda(w_{j}-w_{i}-1)>w_{i}+n^{1/2}$, so $(w_{j}-1)\frac{\lambda+w_{i}}{\lambda+w_{j}}-w_{i}\geq\Omega(n^{-1/2})$ and so the expression displayed above is positive, a contradiction. ∎
@@
$\geq 3\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)-2+\frac{2e(W_{1})}{\lambda}\left(1-\frac{w_{i}}{\lambda+w_{i}}\right)+O(n^{-1})$
@@
Now using $e(W_{1})\leq\frac{n}{r}+O(1)$ we obtain $\frac{2e(W_{1})}{\lambda}\left(\frac{w_{i}}{\lambda+w_{i}}-1\right)\geq-\frac{2}{r}+o(1)$, so
$00\geq 3\left(1-\frac{1}{r}\right)-1-\frac{2}{r}+o(1)\geq 2-\frac{5}{3}+o(1)>$ 0
@@
Let $x$ be the principal eigenvector of $H$, scaled so that its largest entry is $1$. From the estimates proved above, we already have that $x_{v}=1+O(n^{-1})$ for every $v$. Let $A^{\prime}$ be the set of deleted edges and $B^{\prime}$ be the set of added edges. Then we have
@@
By symmetry, we also have
$\lambda(H)-\lambda(H^{\prime})\geq\frac{2(a-b)}{n}+O(n^{-2})$
which proves the result. ∎
@@
which gives a contradiction if $r\geq 4$. If $r=3$, we require a more detailed argument. Without loss of generality, $i=2$. Let $G^{\prime}$ be obtained from $G$ by converting every vertex in $W^{m}$ to an $W$-ordinary vertex in the same part, let $a=e(G^{\prime})-e(G)=O(1)$, and let $G_{21}^{\prime}=(G^{\prime})_{21}$. Observe by 41 that $\lambda(G^{\prime})=\lambda(G)+2a/n+O(n^{-2})$ and $\lambda(G_{21})=\lambda(G_{21}^{\prime})-2a/n+O(n^{-2})$, and so it suffices to show that $\lambda(G_{21}^{\prime})-\lambda(G^{\prime})\geq\delta/n$ for some constant $\delta>0$. Let $V(G_{21}^{\prime})$ have partition $U_{1}\sqcup U_{2}\sqcup U_{3}$ so that $|U_{1}|=w_{1}+1$, $|U_{2}|=w_{2}-1$, and $|U_{3}|=w_{3}$. Note that according to the previous lemmas, we either have $w_{1}=w_{2}-2=w_{3}-1$ or $w_{1}=w_{2}-2=w_{3}-2$. Let $\mu=\lambda(G^{\prime})$, $\rho=\lambda(G_{21}^{\prime})$, with corresponding eigenvectors $y,z$, respectively, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}x_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in U_{1}}z_{v}=1$. Let $y_{2},y_{3}$ be the common eigenweights of vertices in $W_{2},W_{3}$, and let $z_{2},z_{3}$ be the common eigenweights of vertices in $U_{2},U_{3}$.
@@
$\frac{1}{2}(\lambda(G_{21}^{\prime})-\lambda(G^{\prime}))\geq\frac{-w_{1}(1+O(n^{-1}))^{2}+w_{2}(1+O(n^{-1}))^{2}}{y^{T}y}=\frac{O(1)}{y^{T}y}=O(n^{-1}).$
@@
and for the vertex $v\in U_{1}-W_{1}$ we have $z_{v}=1$. Let $Y_{1}=\sum_{v\in W_{1}}y_{v}$, $Z_{1}=\sum_{v\in U_{1}}z_{v}$. Then using the eigenweight estimates above we obtain
@@
Suppose $G_{1}$, $G_{2}$ are graphs obtained from $T_{n,r}$ by embedding the same number of trees of order $k\geq 4$ in a large part so as to cover all but $O(1)$ vertices in the part. Then
@@
Let $G_{\mathrm{star}}$ and $G_{\mathrm{path}}$ be the graphs obtained from $G_{1}$ or $G_{2}$ by replacing each embedded tree with a star or path, respectively.